OpenStudy (kpop4life123):

I WILL FAN AND MEDAL!!!!!!!!!!!!!! What is p^m ÷ p^n equal to? A) p^m – n B) p^m + n C) p^m • n D) p^m ÷ n

OpenStudy (campbell_st):

the law for dividing the same base is \[x^a \div x^b = x^{a - b}~~~or~~~\frac{x^a}{x^b} = x^{a - b}\]

OpenStudy (fibonaccichick666):

so, when multiplying like bases, what do you do to the exponents?

Directrix (directrix):

Follow @campbell_s 's suggestion.

Directrix (directrix):

@Kpop4life123

Directrix (directrix):

Any idea?

OpenStudy (empty):

You shouldn't memorize this so-called "law". It's not a law. It's just a natural consequence of extending the notation. All exponents represent is a smaller way to write multiplication of a number by itself. Once you get used to this, then you see that it naturally extends to negative and fractional exponents.

OpenStudy (empty):

\[x^3*x^2=(x*x*x)*(x*x)=x^5\] Oh the rule is just adding, \[x^nx^m=x^{n+m}\] what if we have negatives? Play with it.

OpenStudy (fibonaccichick666):

that is true, empty, we can prove it quite easily. :) But for some people, it is just easier to memorize things(especially if math is not their strong suit)

OpenStudy (empty):

How do you know that rote memorization is best for this person if they have never even been shown that there's a way to figure it out simply and understand it?

OpenStudy (kpop4life123):

p^(m-n)?

Directrix (directrix):

@Kpop4life123 Correct according to Campbell's exponent rule.

OpenStudy (fibonaccichick666):

well it also happens to be given on most formula sheets in the US. Empty, but I was just answering why someone memorizes the formula. Anyways, wanna prove it for kicks and giggles?

OpenStudy (fibonaccichick666):

also, if people don't get the concept of multiplication just being shorthand for addition-which many many many unfortunately never do- the whole exponents thing is incomprehensible.

OpenStudy (kpop4life123):

why do u make stuff so difficult

OpenStudy (fibonaccichick666):

That was not making it difficult. I was just going to have you relate the skill for multiplying to deduce what the rule is for dividing.